How much 10% solution and how much 45% solution should be mixed together to make 100 gallons of 25% solution?
Answer by jorel1380(3719)

100 x.25=25 gallons extract
.1m+.45(100-m)=25
.1m+45-.45m=25
20=.35m
20/.35=m
m=57.143
100-m=42.857
You need 57.143 gals of 10% solution and 42.857 gals of 45% solution for your mix..